# The Low-Temperature Phenomenology of Gap Inhomogeneity in the Cuprate Superconductors: High-Energy Granularity, Low-Energy Homogeneity, and Spectral Kinks

### v1

@@ II.1 Mean-Field Hamiltonian and Green’s Functions
[…]
@@
$t_{ij}$ and $\Delta_{ij}$ are the hopping matrix element and the bond superconducting order parameter, respectively, between sites $i$ and $j$. The notation $\langle ij\rangle$ in the sum denotes a restriction to only pairs of sites $i$ and $j$ that are nearest-neighbors, next-nearest neighbors, or the same as one another. For $t_{ij}$, we disregard higher-order hoppings, so that the normal-state disperson is $\epsilon_{\mathbf{k}}=-2t(\cos k_{x}+\cos k_{y})-4t^{\prime}\cos k_{x}\cos k_{y}-\mu$, where $t$, $t^{\prime}$, and $\mu$ are the nearest-neighbor hopping, next-nearest-neighbor hopping, and the on-site chemical potential, respectively. We set these hopping parameters to be $t=1$, $t^{\prime}=-0.3$, and $\mu=-1.03$; the resulting hole doping is $16\%$, in line with the real-world hole doping at which the cuprates generally attain the highest $T_{c}$. Meanwhile, for $\Delta_{ij}$, we consider only nearest-neighbor pairs of sites. For a clean $d$-wave superconductor, $\Delta_{\mathbf{r},\mathbf{r}\pm\hat{x}}=\Delta_{0}$ and $\Delta_{\mathbf{r},\mathbf{r}\pm\hat{y}}=-\Delta_{0}$ so that the order parameter in $\mathbf{k}$-space is given by $\Delta_{\mathbf{k}}=2\Delta_{0}(\cos k_{x}-\cos k_{y})$. We will consider in this paper the more general case where $\Delta_{ij}$ is spatially disordered.
@@
We emphasize once more that we keep the hoppings and the on-site chemical potential constant throughout the entire system, so that the standard assumptions of “dirty $d$-wave” theory (where disorder is present through spatial fluctuations of the on-site chemical potential) do not apply here. Instead, inhomogeneity will enter only through the “off-diagonal” order-parameter matrix elements $\Delta_{ij}$. It is worth mentioning that if we had chemical-potential or hopping disorder, it will naturally generate off-diagonal disorder through the self-consistent solution of the BdG Hamiltonian Atkinson et al. (2000a, b); Lee et al. (2016); Sulangi et al. (2021); Li et al. (2021); Pal et al. (2023). However, it is generally found that with weak-to-intermediate levels of chemical-potential disorder (necessary to preserve the observed homogeneity of the superconducting state), the resulting order-parameter inhomogeneity is far too narrowly distributed to be consistent with experimentally obtained gap maps Sulangi et al. (2021). This points to the necessity of including intrinsic disorder in the off-diagonal order-parameter channel that is independent of the variation produced by impurities via self-consistency. On the other hand, a model of glassy nematicity (but with a spatially constant pairing interaction) is able to produce order-parameter variations that look similar to what is seen in experiment Lee et al. (2016). In this work, we do not self-consistently determine the order parameter; instead, we impose a fixed order-parameter profile which is characterized by a particular disorder length scale and strength. A fully self-consistent treatment study of pairing-interaction inhomogeneity and its effects on the superfluid density, the order parameter, and the LDOS at various temperatures will be the subject of a forthcoming paper Sulangi et al. .
@@
Here, the $\mathbf{M}_{i}$ and $\mathbf{T}_{i}$ blocks are $2N_{y}\times 2N_{y}$ matrices which encode the couplings for both particle and hole degrees of freedom. $\mathbf{M}_{i}$ contains all couplings among sites belonging to the $i$th column, while $\mathbf{T}_{i}$ couples sites in the $i$th column with those in the $i+1$th column. One can thus view the system as quasi-one dimensional in the $x$-direction, with $\mathbf{T}_{i}$ serving as the “hopping” and $\mathbf{M}_{i}$ as the “on-site energy.” The diagonal blocks of $\mathbf{G}$ can then be obtained by a recursive algorithm with computational complexity $\mathcal{O}(N_{x}N_{y}^{3})$ for each frequency at which the calculation is performed Godfrin (1991); Reuter and Hill (2012). To be more specific, we introduce two sets of auxilliary matrices $\mathbf{A}_{i}$ and $\mathbf{B}_{i}$, all with dimension $2N_{y}\times 2N_{y}$ and defined recursively in the following way:
$\mathbf{A}_{i}=\begin{cases}\mathbf{0}&\text{for }i=1,\\ \mathbf{T}_{i-1}^{\dagger}\Bigl{[}(\omega+i\Gamma)\mathbf{1}-\mathbf{M}_{i-1}-\mathbf{A}_{i-1}\Bigr{]}\mathbf{T}_{i-1}&\text{for }1<i\leq N_{x}\\ \end{cases}$ (4)
@@
$\mathbf{B}_{i}=\begin{cases}\mathbf{0}&\text{for }i=N_{x},\\ \mathbf{T}_{i}\Bigl{[}(\omega+i\Gamma)\mathbf{1}-\mathbf{M}_{i+1}-\mathbf{B}_{i+1}\Bigr{]}\mathbf{T}_{i}^{\dagger}&\text{for }1\leq i<N_{x}.\\ \end{cases}$ (5)
@@ II.2 Order-Parameter Inhomogeneity
The main objective of this paper is to examine the role of the length scale of order-parameter inhomogeneity on the LDOS and the various observables that can be extracted from it. It is important then to pick a model whose length scale is easily tuned, keeping everything else the same. There is no lack of models of order-parameter inhomogeneity which have appeared in the literature. One extensively studied model is the Yukawa-like model used by Nunner et al., in which off-plane impurities give rise to a moduled pairing interaction whose distance-dependence is that of a screened Coulomb potential Nunner et al. (2005, 2006); Andersen et al. (2006). While realistic in that one can use it to reproduce the gap distributions seen in experiment, the overall inhomogeneity length scale in this model is not easily tuned. Three length scales are present in the Yukawa problem: the distance between the impurity and the CuO2 plane, the screening length, and the mean distance between impurities. The resulting overall length scale $l$ of the order parameter inhomogeneity then has a nontrivial dependence on the interplay among these three Yukawa length scales, making it not straightforward to adjust $l$.
Instead, we use as a basis the simpler model first studied by Fang et al. Fang et al. (2006). The model consists of a square patch of length $l$ within which the $d$-wave order parameter takes on a constant value, motivated by STS gap maps that show a patchy structure of the gap with what appears to be “phase separation” between small-gap and large-gap regions. (A similar square-patch model was used by Alvarez and Dagotto, but with random phases instead of amplitudes on each patch Alvarez and Dagotto (2008).) Instead of considering a single isolated patch as Fang et al. did, we instead consider a “many-patch” version of this disorder model. The entire $N_{x}\times N_{y}$ system is subdivided into square regions, each with size $l\times l$. Let us first define $\Delta_{D,i}$ as the site-centered order parameter at site $i$ from which the bond order parameters (which enter the Hamiltonian) are derived. For each square patch, $\Delta_{D,i}$ is constant for all sites $i$ within that patch. The bond order parameter $\Delta_{ij}$ (which enters Eq. 1) is then obtained from $\Delta_{D,i}$ via the following relation:
@@
The values of $\Delta_{D}$ on each patch are independent and identically distributed, with their values taken from a uniform distribution with mean $\overline{\Delta}$ and width $2\overline{\Delta}$. This choice of distribution ensures that a wide range of order parameter values is accessed. In particular, both large-gap patches (with $\Delta_{D}\approx 2\overline{\Delta}$) and small-gap patches (with $\Delta_{D}\approx 0$) coexist here. While our model for gap inhomogeneity does not appear to be similar to known gap distributions from STS experiments, it bears emphasizing that the gap extracted from experiment is not exactly the same as the underlying order parameter, and we make no attempt at fine-tuning the parameters to ensure that this model accurately reproduces the experimentally obtained distributions. Nevertheless, we will show that in certain regimes, the resulting distribution of the gaps becomes similar to that seen in experiment without any fine-tuning, the square shape of the underlying patches notwithstanding.
@@
We will also consider the maximally disordered case where the order parameter on every bond is independent and identically distributed from a uniform distribution with mean $\frac{\overline{\Delta}}{4}$ and width $\frac{\overline{\Delta}}{2}$ (the $\frac{1}{4}$ prefactor again is needed for proper normalization of the bond order parameter). We will call this the “$l=0$” case as it is the extreme limit where no spatial correlations are present between the bond order parameters. Note that the smallest patch in our disorder model ($l=1$) still exhibits correlations among adjacent bond order parameters, as these are derived from $\Delta_{D}$ via Eq. 8.
@@
Because the objective of this paper is to examine the effect of the length scales of the patchy inhomogeneity, we will neglect the addition of oscillatory components of the $d$-wave order parameter that have been identified in recent STS experiments on Bi-2212 Hamidian et al. (2016); Edkins et al. (2019); Du et al. (2020) and Bi-2223 Zou et al. (2022). This oscillatory component, which is found to be much smaller relative to the uniform component, has been identified as a “pair-density wave” in Bi-2212, but an alternative disorder-based mechanism involving quasiparticle scattering between antinodal states has been proposed for this which accounts for the particle-hole asymmetry of the oscillations in the coherence-peak location Zou et al. (2022); Gao et al. (2024). We will not include these ingredients into the present paper, and only focus on the inhomogeneity of the background (i.e., non-oscillatory) $d$-wave order parameter.
@@
The only parameters that we tune here are the size $l$ of the patches and the mean $\overline{\Delta}$; the latter sets the average coherence length $\xi$, and how large $l$ is compared to $\xi$ will turn out to be of significance to many observables. We will show results primarily for $\overline{\Delta}=0.16$ (with a coherence length $\xi\approx 4.53$ lattice spacings, this is a fairly accurate representation of mildly overdoped superconducting cuprates), but we will also present calculations for $\overline{\Delta}=0.32$ ($\xi\approx 2.27$) and $\overline{\Delta}=0.08$ ($\xi\approx 9.06$).
II.3 Order Parameter Proxies

### v2

@@ II.1 Mean-Field Hamiltonian and Green’s Functions
$H=\sum_{\langle ij\rangle}\Big[-\sum_{\sigma}t_{ij}c_{i\sigma}^{\dagger}c_{j\sigma}+\Delta_{ij}^{\ast}c_{i\uparrow}c_{j\downarrow}+\text{h.c.}\Big].$ (1)
@@
$t_{ij}$ and $\Delta_{ij}$ are the hopping matrix element and the bond superconducting order parameter, respectively, between sites $i$ and $j$. The notation $\langle ij\rangle$ in the sum denotes a restriction to only pairs of sites $i$ and $j$ that are nearest-neighbors, next-nearest neighbors, or the same as one another. For $t_{ij}$, we disregard higher-order hoppings, so that the normal-state dispersion is $\epsilon_{\mathbf{k}}=-2t(\cos k_{x}+\cos k_{y})-4t^{\prime}\cos k_{x}\cos k_{y}-\mu$, where $t$, $t^{\prime}$, and $\mu$ are the nearest-neighbor hopping, next-nearest-neighbor hopping, and the on-site chemical potential, respectively. We set these hopping parameters to be $t=1$, $t^{\prime}=-0.3$, and $\mu=-1.03$; the resulting hole doping is $16\%$, in line with the real-world hole doping at which the cuprates generally attain the highest $T_{c}$. Meanwhile, for $\Delta_{ij}$, we consider only nearest-neighbor pairs of sites. For a clean $d$-wave superconductor, $\Delta_{\mathbf{r},\mathbf{r}\pm\hat{x}}=\Delta_{0}$ and $\Delta_{\mathbf{r},\mathbf{r}\pm\hat{y}}=-\Delta_{0}$ so that the order parameter in $\mathbf{k}$-space is given by $\Delta_{\mathbf{k}}=2\Delta_{0}(\cos k_{x}-\cos k_{y})$. We consider in this paper the more general case where $\Delta_{ij}$ is spatially disordered.
@@
We emphasize once more that we keep the hoppings and the on-site chemical potential constant throughout the entire system, so that the standard assumptions of “dirty $d$-wave” theory (where disorder is present through spatial fluctuations of the on-site chemical potential) do not apply here. Instead, inhomogeneity will enter only through the “off-diagonal” order-parameter matrix elements $\Delta_{ij}$. It is worth mentioning that if we had chemical-potential or hopping disorder, it will naturally generate off-diagonal disorder through the self-consistent solution of the BdG Hamiltonian [42, 43, 37, 44, 45, 46]. However, it is generally found that with weak-to-intermediate levels of chemical-potential disorder (necessary to preserve the observed homogeneity of the superconducting state), the resulting order-parameter inhomogeneity is far too narrowly distributed to be consistent with experimentally obtained gap maps [44]. This points to the necessity of including intrinsic disorder in the off-diagonal order-parameter channel that is independent of the variation produced by impurities via self-consistency. On the other hand, a model of glassy nematicity (but with a spatially constant pairing interaction) is able to produce order-parameter variations that look similar to what is seen in experiment [37]. In this work, we do not self-consistently determine the order parameter; instead, we impose a fixed order-parameter profile which is characterized by a particular disorder length scale and strength. A fully self-consistent treatment study of pairing-interaction inhomogeneity and its effects on the superfluid density, the order parameter, and the LDOS at various temperatures will be the subject of a forthcoming paper [47].
@@
Here, the $\mathbf{M}_{i}$ and $\mathbf{T}_{i}$ blocks are $2N_{y}\times 2N_{y}$ matrices which encode the couplings for both particle and hole degrees of freedom. $\mathbf{M}_{i}$ contains all couplings among sites belonging to the $i$th column, while $\mathbf{T}_{i}$ couples sites in the $i$th column with those in the $i+1$th column. One can thus view the system as quasi-one dimensional in the $x$-direction, with $\mathbf{T}_{i}$ serving as the “hopping” and $\mathbf{M}_{i}$ as the “on-site energy.” The diagonal blocks of $\mathbf{G}$ can then be obtained by a recursive algorithm with computational complexity $O(N_{x}N_{y}^{3})$ for each frequency at which the calculation is performed [48, 49]. To be more specific, we introduce two sets of auxilliary matrices $\mathbf{A}_{i}$ and $\mathbf{B}_{i}$, all with dimension $2N_{y}\times 2N_{y}$ and defined recursively in the following way:
[…]
@@
[…]
@@ II.2 Order-Parameter Inhomogeneity
The main objective of this paper is to examine the role of the length scale of order-parameter inhomogeneity on the LDOS and the various observables that can be extracted from it. It is important then to pick a model whose length scale is easily tuned, keeping everything else the same. There is no lack of models of order-parameter inhomogeneity which have appeared in the literature. One extensively studied model is the Yukawa-like model used by Nunner et al., in which off-plane impurities give rise to a modulated pairing interaction whose distance-dependence is that of a screened Coulomb potential [39, 50, 40, 41]. While realistic in that one can use it to reproduce the gap distributions seen in experiment, the overall inhomogeneity length scale in this model is not easily tuned. Three length scales are present in the Yukawa problem: the distance between the impurity and the CuO2 plane, the screening length, and the mean distance between impurities. The resulting overall length scale $l$ of the order parameter inhomogeneity then has a nontrivial dependence on the interplay among these three Yukawa length scales, making it not straightforward to adjust $l$.
Instead, we use as a basis the simpler model first studied by Fang et al. [8]. The model consists of a square patch of length $l$ within which the $d$-wave order parameter takes on a constant value, motivated by STS gap maps that show a patchy structure of the gap with what appears to be “phase separation” between small-gap and large-gap regions. (A similar square-patch model was used by Alvarez and Dagotto, but with random phases instead of amplitudes on each patch [38].) Instead of considering a single isolated patch as Fang et al. did, we instead consider a “many-patch” version of this disorder model. The entire $N_{x}\times N_{y}$ system is subdivided into square regions, each with size $l\times l$. Let us first define $\Delta_{D,i}$ as the site-centered order parameter at site $i$ from which the bond order parameters (which enter the Hamiltonian) are derived. For each square patch, $\Delta_{D,i}$ is constant for all sites $i$ within that patch. The bond order parameter $\Delta_{ij}$ (which enters Eq. 1) is then obtained from $\Delta_{D,i}$ via the following relation:
@@
The values of $\Delta_{D}$ on each patch are independent and identically distributed, with their values taken from a uniform distribution with mean $\overline{\Delta}$ and width $2\overline{\Delta}$. This choice of distribution ensures that a wide range of order parameter values is accessed. In particular, both large-gap patches (with $\Delta_{D}\approx 2\overline{\Delta}$) and small-gap patches (with $\Delta_{D}\approx 0$) coexist here. While our model for gap inhomogeneity does not appear to be similar to known gap distributions from STS experiments, it bears emphasizing that the gap extracted from experiment is not exactly the same as the underlying order parameter, and we make no attempt at fine-tuning the parameters to ensure that this model accurately reproduces the experimentally obtained distributions. Nevertheless, we show that in certain regimes, the resulting distribution of the gaps becomes similar to that seen in experiment without any fine-tuning, the square shape of the underlying patches notwithstanding.
@@
We also consider the maximally disordered case where the order parameter on every bond is independent and identically distributed from a uniform distribution with mean $\frac{\overline{\Delta}}{4}$ and width $\frac{\overline{\Delta}}{2}$ (the $\frac{1}{4}$ prefactor again is needed for proper normalization of the bond order parameter). We call this the “l=0l=0” case as it is the extreme limit where no spatial correlations are present between the bond order parameters. Note that the smallest patch in our disorder model ($l=1$) still exhibits correlations among adjacent bond order parameters, as these are derived from $\Delta_{D}$ via Eq. 8.
@@
Because the objective of this paper is to examine the effect of the length scales of the patchy inhomogeneity, we neglect the addition of oscillatory components of the $d$-wave order parameter that have been identified in recent STS experiments on Bi-2212 [20, 21, 22] and Bi-2223 [51]. This oscillatory component, which is found to be much smaller relative to the uniform component, has been identified as a “pair-density wave” in Bi-2212, but an alternative disorder-based mechanism involving quasiparticle scattering between antinodal states has been proposed for this which accounts for the particle-hole asymmetry of the oscillations in the coherence-peak location [51, 52]. We will not include these ingredients into the present paper, and only focus on the inhomogeneity of the background (i.e., non-oscillatory) $d$-wave order parameter.
@@
The only parameters that we tune here are the size $l$ of the patches and the mean $\overline{\Delta}$; the latter sets the average coherence length $\xi$, and how large $l$ is compared with $\xi$ will turn out to be of significance to many observables. We show results primarily for $\overline{\Delta}=0.16$ (with a coherence length $\xi\approx 4.53$ lattice spacings, this is a fairly accurate representation of mildly overdoped superconducting cuprates), but we also present calculations for $\overline{\Delta}=0.32$ ($\xi\approx 2.27$) and $\overline{\Delta}=0.08$ ($\xi\approx 9.06$).
II.3 Microscopic Justification for the Order-Parameter Inhomogeneity Model
Even though our model appears to be artificial as it involves square patches, it is intended to capture important aspects of the cuprates within the simplest model of inhomogeneity possible. We highlight two features of this model that (with the aid of hindsight) are central to the phenomenology of the cuprates. The first aspect is the presence of a length scale (in our model, the size of the patches $l$) that governs the range of spatial correlations of the order parameter inhomogeneity. The second aspect is an extremely broad distribution of order-parameter values. In our treatment, both the inhomogeneity length scale and the width of the disorder are tunable parameters that we do not tie to a specific microscopic model. One can ask if these phenomenological assumptions are well-founded. In this section, we discuss a specific model that can give rise to the salient features of our chosen order-parameter inhomogeneity model, and discuss two possible microscopic realizations of this in the literature.
We first note that one can obtain these order-parameter distributions by assuming the existence of a spatially varying attractive nearest-neighbor pairing interaction $V_{ij}$ (here, $i$ and $j$ label sites on the square lattice). In a self-consistent mean-field treatment, the bond order parameter $\Delta_{ij}$ is given by [53]
$\Delta_{ij}=V_{ij}\langle c_{i\uparrow}c_{j\downarrow}\rangle.$ (9)
With a suitably chosen spatial profile of the pairing interaction $V_{ij}$, one can in principle reproduce any order-parameter inhomogeneity profile, including ours. We in fact show this in detail in a detailed follow-up work [47], where we demonstrate that broadly distributed square-patch inhomogeneity can arise within self-consistent mean-field theory, assuming that the pairing interaction $V_{ij}$ itself is also broadly distributed and follows the square-patch inhomogeneity model. It is evident that a microscopic model underpinning our phenomenological one involves a spatially varying pairing interaction treated self-consistently [39, 50, 8, 41, 54, 16, 53]. Further aspects of pairing-interaction inhomogeneity, including its effects on $T_{c}$ and the superfluid density, have previously been discussed in Refs. [55, 56, 57] and will be expounded upon further in our upcoming work [47].
It is helpful to recap why inhomogeneity in the pairing interaction is likely to be present in the cuprates, and why this is our preferred starting point in our analysis. Nunner et al. find that many correlations seen among STS observables (e.g., the local spectral gap, the coherence-peak height, and dopant locations), as well as particle-hole symmetric local spectra, are reproduced nicely within mean-field theory if one assumes that the dopant atoms cause local modulations in the pairing interaction [39, 50]. Similarly, Fang et al. also find that the anticorrelation between the spectral gap size and the coherence peak height arises naturally from nanoscale inhomogeneity in the order parameter alone [8]. In addition, Sulangi et al. show that a good deal of the cuprate STS phenomenology is reproduced using the pairing-interaction inhomogeneity model of Nunner et al. treated within both self-consistent mean-field theory and time-dependent Ginzburg-Landau theory, with the latter demonstrating consistency with STS experiments done at high temperatures [53].
It is important to note that if one instead assumes a constant pairing interaction and inhomogeneity in the chemical potential (with order-parameter inhomogeneity generated self-consistently), many of the aforementioned features are not easily reproduced within mean-field theory. In particular, Nunner et al. [39, 50] and Sulangi et al. [44] find that within such a scenario, the LDOS near impurities are quite particle-hole asymmetric, in contrast with experiment. More importantly, for pure chemical-potential disorder to generate the order-parameter variations seen in experiment, this has to be strong enough that large local charge modulations [39], subgap states, and gap filling [46] are generated, which are not seen in STS. As gap filling is absent at low temperatures in the cuprates throughout most of the doping ranges we are concerned with, it is highly unlikely then that the inhomogeneity in the order parameter is the result of mean-field backreaction of the order parameter on strong on-site potentials. No evidence for kinks and a low-energy–high-energy dichotomy akin to experiment is seen in self-consistent mean-field models involving only on-site potential disorder [46].
Inhomogeneity in the hopping amplitudes has been considered by Lee et al. [37]. Here, they assume a uniform pairing interaction, but the hoppings are modulated (with some correlation length) in such a way that a “nematic glass” is formed; their treatment is entirely mean-field in character. The generated superconducting order parameter is spatially extended, with the length scale set by that of the underlying nematic inhomogeneity. It is not clear from their analysis if the correlations among STS observables are reproduced within their model. However, they find a version of the low-energy–high-energy dichotomy in their calculated LDOS spectra. However, they do not appear to be able to reproduce kinks in the LDOS within their model. Instead, they find that the LDOS is homogeneous at low energies below the average gap energy, and inhomogeneous above it, a result not similar to what is seen in STS experiments. It is not known if this discrepancy is because the resulting order-parameter variation is too narrow (compare with our results in Appendix A for some results on the order parameter distribution width-dependence of the LDOS), or if this is due to effects peculiar to the normal-state properties of their model of glassy nematicity.
What can cause an inhomogeneous pairing interaction to form? It is likely that this is due to the correlated nature of the cuprates. This is the point of view that Rømer et al. [58, 54] and Chakraborty et al. [59, 60, 61] adopt. Rømer et al. [58, 54] make use of a real-space formulation of spin fluctuations (treated within the random-phase approximation) to derive an effective pairing interaction, which owing to self-consistency will be sensitive to the presence of disorder in the on-site potential and in the hoppings (in their paper, they only consider site disorder in the form of nonmagnetic impurities). They then feed this inhomogeneous pairing interaction into a self-consistent Bogoliubov-de Gennes calculation, yielding an inhomogeneous $d$-wave superconductor. The resulting spectral gap maps with randomly distributed impurities are strikingly patchy and are reminiscent of experiment. This work represents a proof of principle that enhanced pairing can be “nucleated” in the vicinity of pointlike impurities, thereby providing a concrete microscopic model underlying the one used in Nunner et al. In particular, the enhancement of the spectral gap was found to extend several lattice spacings around a single isolated impurity, and particularly strong enhancements in turn can be generated by rare diagonal “dimer” impurity configurations. These allow for the generation of broadly distributed and spatially extended order parameter inhomogeneity; the effects are even more pronounced the closer one gets to the magnetic instability point. A similar observation is made by a first-principles study which finds that the pairing interaction is enhanced in the vicinity of a oxygen dopant in Bi-2212 [62].
Chakraborty et al. [59, 60, 61] on the other hand obtain a Gutzwiller-renormalized normal-state Hamiltonian. The Gutzwiller factors that encode the effects of strong correlations then enter into the self-consistent expression for the superconducting order parameter; a spatially dependent pairing interaction is generated as a result. It is worth noting that in the correlated calculation, they find the inhomogeneity in the resulting order parameter to be less sensitive to the underlying disorder than in the uncorrelated case. It would appear that within the Gutzwiller-renormalized theory, much stronger disorder needs to be present to reproduce the broad gap distributions seen in experiment [60]. Correlation effects as modeled by the Gutzwiller-renormalized theory tend to keep the LDOS surprisingly uniform even in the presence of strong disorder, but this may partially be due to the small length scale (on the order of the Fermi wavelength $k_{F}^{-1}$) of the ensuing order-parameter inhomogeneity, which would give rise to a largely uniform quasiparticle response. (For comparison and corroborating evidence from a mean-field calculation, please see our $l<\xi$ results in the next section.)
In any case, the differences between Ref. [54] and Refs. [59, 60, 61] highlight the sensitivity of the resulting observables to the particulars of the correlation model used. This is especially so when the ultimate source of inhomogeneity is not settled with any sort of definiteness. We have therefore opted to treat inhomogeneity entirely within a phenomenological model, and we have not included correlation effects in the normal-state Hamiltonian, similar to Nunner et al. [39, 50], Fang et al. [8], and Tromp et al. [16]. This “back-to-basics” correlation-agnostic approach ultimately allows us to isolate what non-correlation-related ingredients are necessary for the phenomenology of STS experiments on superconducting cuprates across a wide doping range, over which the strength of the electron-electron correlations presumably varies considerably.
II.4 Order Parameter Proxies